# Confidence as Forecast: A Decision-Theoretic Interpretation of Confidence Intervals

### v1

@@ 5.2 Isn’t this epistemic?
Part of why Neyman may have resisted a forecasting-based interpretation of ex-post coverage is that it threatens to blur the boundary between his frequentist, procedural, behavioristic view of inference and the then-emerging subjectivist Bayesian program of Ramsey [16] and de Finetti [2]. In particular, the gap between Neyman’s error-controlling constructions and de Finetti’s slogan that “probability does not exist” (ibid.) was likely a bridge too far. Still, as the preceding sections suggest, when we treat “confidence” as a predictive probability for an underlying coverage event determined by the design, the resulting notion remains both objective and purely frequentist. Both ex ante and ex post, coverage probabilities are defined as limiting relative frequencies of success in clearly specified reference classes, e.g., intervals produced by this procedure, or intervals of a given relative width $W$ under a given design, and under any strictly proper scoring rule, the risk-minimizing forecast is exactly the relevant conditional coverage probability, whether unconditional ($1-\alpha$) or conditioned on a $\theta$-free statistic such as relative width. So long as those conditional probabilities are all equally well-defined under the sampling model, there seems to be no mathematical reason to treat some as more legitimate than others, or to insist that our confidence in coverage must not change, even if the design itself tells us how it should.
@@
Taking a slightly more philosophical tack, we might also note that Neyman’s behavioristic program is explicitly designed to control our long-run error rates in making statistical inferences, for example in stating that an interval covers the parameter, or in rejecting a null hypothesis. The only reason we ever make such errors, however, is that we, as agents, do not know whether the claims themselves are true. If we did—as in the familiar stacked-interval plot where $\theta$ is taken to be known, and we can see exactly which intervals cover—our error rate would be zero, and there would be no substantive need for confidence procedures in the first place. It is precisely our lack of knowledge that makes long-run error control, and hence both Neyman’s framework and the error-statistical machinery built more recently around it [10], both useful and relevant. To be sure, though, none of this is to suggest that CI theory itself is subjective in any way, or that the conditional coverage probabilities explored in Sections 3 and 4 are somehow tied to degree of belief, credence, propensity, or any other philosophical interpretation of probability aside from frequentism (again, they are defined purely with respect to limit relative frequencies under repeated sampling). Mainly, it is to suggest that the idea of ”confidence”, as Neyman conceived, was defined relative to an agent’s (typically a statistician’s) information state when making statements about unobserved events, and could thus be fairly deemed to be at least information-relative, if not actually epistemic in the subjective sense.

### v2

@@ 5.2 Isn’t this epistemic?
Part of why Neyman may have resisted a forecasting-based interpretation of ex-post coverage is that it threatens to blur the boundary between his frequentist, procedural, behavioristic view of inference and the then-emerging subjectivist Bayesian program of Ramsey [16] and de Finetti [2]. In particular, the gap between Neyman’s error-controlling constructions and de Finetti’s slogan that ’probability does not exist’ (ibid.) may have represented a sharper break than Neyman was willing to make. Still, as the preceding sections suggest, when we treat “confidence” as a predictive probability for an underlying coverage event determined by the design, the resulting notion remains both objective and purely frequentist. Both ex ante and ex post, coverage probabilities are defined as limiting relative frequencies of success in clearly specified reference classes, e.g., intervals produced by this procedure, or intervals of a given relative width $W$ under a given design, and under any strictly proper scoring rule, the risk-minimizing forecast is exactly the relevant conditional coverage probability, whether unconditional ($1-\alpha$) or conditioned on a $\theta$-free statistic such as relative width. So long as those conditional probabilities are all equally well-defined under the sampling model, there seems to be no mathematical reason to treat some as more legitimate than others, or to insist that our confidence in coverage must not change, even if the design itself tells us how it should.
@@
Taking a slightly more philosophical tack, we might also note that Neyman’s behavioristic program is explicitly designed to control our long-run error rates in making statistical inferences, for example in stating that an interval covers the parameter, or in rejecting a null hypothesis. The only reason we ever make such errors, however, is that we, as agents, do not know whether the claims themselves are true. If we did—as in the familiar stacked-interval plot where $\theta$ is taken to be known, and we can see exactly which intervals cover—our error rate would be zero, and there would be no substantive need for confidence procedures in the first place. It is precisely our lack of knowledge that makes long-run error control, and hence both Neyman’s framework and the error-statistical machinery built more recently around it [10], both useful and relevant. To be sure, though, none of this is to suggest that CI theory itself is subjective in any way, or that the conditional coverage probabilities explored in Sections 3 and 4 are somehow tied to degree of belief, credence, propensity, or any other philosophical interpretation of probability aside from frequentism (again, they are defined purely with respect to limit relative frequencies under repeated sampling). Mainly, it is to suggest that the idea of ”confidence”, as Neyman conceived, was defined relative to an agent’s (typically a statistician’s) information state when making statements about unobserved events, and could thus reasonably be described as at least information-relative, if not actually epistemic in the subjective sense.